A space station shaped like a giant wheel has a ra
Answer #1


Ii = 4.96*10^8 kg.m^2, r =109 m , ai =1g,

angular acceleration αi = a/r =9.8/r

from coservationtion of momentum

Iiαi = Ifαf

If = Ii - m*150*r^2

(4.96*10^8*9.80/r = (4.96*10^8 - (65*150*109*109))(af/r)

af =12.79 m/s^2